# Cinematic Gaussians: Real-Time HDR Radiance Fields with Depth of Field

### v2

@@ 4.2.2 HDR Reconstruction
For HDR reconstruction, we compare our results against HDR-NeRF [HZF22] using various metrics specifically targeted towards evaluating HDR quality, including HDR-VDP-3 [MHH23], PU-PSNR, and PU-SSIM [A21], as summarized in Table 2. Visual comparisons with the ground truth are depicted in Fig. 4. Again, we use for evaluation the four scenes of our synthetic dataset, with 64 (8 $\times$ 8) views for training and 49 (7 $\times$ 7) novel ones for testing. To simulate LDR training views with varying exposures, starting from the HDR renders from our synthetic dataset, we sample a single exposure for each view. For this, we first determine the start and end boundaries that emcompass the entire dynamic range of the scene [ANSAM21]. Then, we divide this range into eight intervals, within which we sample exposures as described in previous work [HZF22]. Our method consistently outperforms HDR-NeRF in almost all metrics, demonstrating superior reconstruction capabilities. Notably, we observed that HDR-NeRF tends to overfit to the LDR images, resulting in lower quality HDR image generation. Moreover, our approach achieves real-time rendering capabilities, significantly outperforming HDR-NeRF. We deliver approximately 110 FPS at a resolution of 1200 x 675, in contrast to HDR-NeRF’s rate of approximately 0.021 FPS for rendering a single image, both using a GPU RTX A40.
Table 2: Comparison of HDR reconstruction performance metrics (HDR-VDP-3, PU-PSNR, and PU-SSIM) for our method and HDR-NeRF using our synthetic dataset.
@@
Figure 4: Comparison of HDR reconstruction. Left-most column presents input LDR images for a given view. The following columns depict results acquired by HDR-NeRF, Ground Truth, and results of our method, respectively. Please note, that our method properly reconstructs the shape of the flame and reproduces successfully high-frequency content, while these details are missing in the results generated by HDR-NeRF. To depict the accuracy of both methods we provide the HDR-VDP-3 error map in the top right corner of the images.
@@ 4.3 HDR Radiance Field Reconstruction and Post-editing
In this section, we showcase the full capabilities of our model, specifically its ability to reconstruct HDR radiance fields, which enable the generation of all-in-focus HDR images. Fig. 5 presents results from various viewpoints, illustrating how our approach effectively reconstructs detailed scenes with consistent sharpness and dynamic range from multi-view LDR defocused inputs. Furthermore, we demonstrate the post-editing capabilities enabled by our method. Our proposed approach reconstructs HDR radiance fields with real-time re-exposure and re-focusing capabilities. Users have the flexibility to adjust aperture size, exposure time, and focal distance in real-time, as demonstrated in Fig. 6.
@@ 4.4 Ablation Study
We conduct this evaluation differently for our synthetic and real datasets. For the synthetic dataset, where ground truth images are available, we evaluate our all-in-focus HDR reconstructions against ground truth views not included in the training set. Specifically, we utilize 64 (8 $\times$ 8) views for training and 49 (7 $\times$ 7) novel views for testing. This comparison is performed in HDR space, therefore we use PU-PSNR, PU-SSIM and HDR-VDP-3 as metrics. We show these results in Table 3. On the other hand, for the captured dataset, where ground truth is not available, we adopt a different approach. We use 74 captured views for training and leave out 7 for testing. Then, we synthesize the testing views with our model using the original aperture, focal distance, and exposure settings corresponding to each captured view, and compare these synthesized views against the actual captures. This comparison is conducted in LDR space, thus we employ traditional image quality metrics, in particular, PSNR, SSIM, and LPIPS. Quantitative results are presented in Table 4, while qualitative results are illustrated in Fig. 7 for the defocus loss term and in Fig. 8 for the exposure loss term.
@@
Methods PU-PSNR($\uparrow$) PU-SSIM($\uparrow$) HDR-VDP($\uparrow$)
Ours 34.04 $\pm$ 1.67 0.95 $\pm$ 0.00 9.83 $\pm$ 0.05
@@
Table 4: Evaluation of the impact of the defocus and exposure loss terms in the captured dataset.
Methods PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Ours 25.69 $\pm$ 2.71 0.86 $\pm$ 0.07 0.10 $\pm$ 0.02
@@ 5 Limitations and Future Work
Our approach does not come without limitations, opening up fruitful avenues for future work.
Our defocus model is a two-fold approximation of the actual physical processes in a real camera lens. First, we use a thin-lens model [PC81], which, despite being commonly used, is a simplified representation of defocus. More sophisticated models have been developed in the forward-rendering literature [KMH95] and related work on defocus modeling in the context of reconstruction from image stacks [WSP23b] has shown that advanced effects, such as lens breathing, can significantly improve accuracy. Second, we approximate the thin-lens circle of confusion with a Gaussian function to achieve real-time rendering performance. While this is a commonly used design decision [Fav10, SZW23], it results in a slight softening of bokeh due to the deviation from a disk kernel. We believe that moving towards more realistic lens models has the potential to further increase result quality and cinematic appeal.
In scenes characterized by extremely high dynamic range, where very dark regions may contain significant noise across most input images, artifacts may occur in the reconstruction of these dark areas. These artifacts are pronounced when selecting very high exposures (Fig. 9). This is expected, since there is insufficient useful information, with noise dominating instead in these areas. Nevertheless, these artifacts are noticeable only under extreme exposure settings in the reconstruction, while our method remains capable of delivering pleasing results for the rest of the scene.

### v3

@@ 4.2.2 HDR Reconstruction
For HDR reconstruction, we compare our results against HDR-NeRF [HZF22] using various metrics specifically targeted towards evaluating HDR quality, including HDR-VDP-3 [MHH23], PU-PSNR, and PU-SSIM [MA21], as summarized in Table 3. Visual comparisons with the ground truth are depicted in Fig. 4. Again, we use for evaluation the four scenes of our synthetic dataset, with 64 (8 $\times$ 8) views for training and 49 (7 $\times$ 7) novel ones for testing. To simulate LDR training views with varying exposures, starting from the HDR renders from our synthetic dataset, we sample a single exposure for each view. For this, we first determine the start and end boundaries that emcompass the entire dynamic range of the scene [ANSAM21]. Then, we divide this range into eight intervals, within which we sample exposures as described in previous work [HZF22]. Our method consistently outperforms HDR-NeRF in almost all metrics, demonstrating superior reconstruction capabilities. Notably, we observed that HDR-NeRF tends to overfit to the LDR images, resulting in lower quality HDR image generation. Moreover, our approach achieves real-time rendering capabilities, significantly outperforming HDR-NeRF. We deliver approximately 110 FPS at a resolution of 1200 x 675, in contrast to HDR-NeRF’s rate of approximately 0.021 FPS for rendering a single image, both using a GPU RTX A40. We also run HDR-NeRF on our real dataset using non-reference metrics PU-PIQE [HME22] to assess results from novel views (Table 4). These results highlight that our method outperforms task-specific approaches.
Table 2: Comparison of all-in-focus reconstruction performance metrics for our method and Deblur- Splatting using our real dataset.
Method NIQE $\downarrow$ BRISQUE $\downarrow$
Deblur-Splatting 6.65 $\pm$ 1.87 45.77 $\pm$ 8.50
Ours 2.66 $\pm$ 0.22 32.17 $\pm$ 4.75
Table 3: Comparison of HDR reconstruction performance metrics for our method and HDR-NeRF using our synthetic dataset.
@@
Table 4: Comparison of HDR reconstruction performance metric for our method and HDR-NeRF using our real dataset.
Method PU-PIQE $\downarrow$
HDR-NeRF 55.47 $\pm$ 2.22
Ours 42.22 $\pm$ 1.22
Figure 4: Comparison of HDR reconstruction. Left-most column presents input LDR images for a given view. The following columns depict results acquired by HDR-NeRF, Ground Truth, and results of our method, respectively. Please note, that our method properly reconstructs the shape of the flame and reproduces successfully high-frequency content, while these details are missing in the results generated by HDR-NeRF. To depict the accuracy of both methods we provide the HDR-VDP-3 error map in the top right corner of the images. To better show the differences in the highlights, we render these regions at -3 stops, as shown in the yellow box.
@@ 4.3 HDR Radiance Field Reconstruction and Post-Editing
In this section, we showcase the full capabilities of our model, specifically its ability to reconstruct HDR radiance fields, which enable the generation of all-in-focus HDR images. Fig. 5 presents results from various viewpoints, illustrating how our approach effectively reconstructs detailed scenes with consistent sharpness and dynamic range from multi-view LDR defocused inputs. Furthermore, we demonstrate the post-editing capabilities enabled by our method. Our proposed approach reconstructs HDR radiance fields with real-time re-exposure and re-focusing capabilities. Users have the flexibility to adjust aperture size, exposure time, and focus distance in real-time, as demonstrated in Fig. 6.
@@ 4.4 Ablation Study
We conduct this evaluation differently for our synthetic and real datasets. For the synthetic dataset, where ground-truth images are available, we evaluate our all-in-focus HDR reconstructions against ground-truth views not included in the training set. Specifically, we utilize 64 (8 $\times$ 8) views for training and 49 (7 $\times$ 7) novel views for testing. This comparison is performed in HDR space, therefore we use PU-PSNR, PU-SSIM and HDR-VDP-3 as metrics. We show these results in Table 5. On the other hand, for the captured dataset, where ground truth is not available, we adopt a different approach. We use 74 captured views for training and leave out 7 for testing. Then, we synthesize the testing views with our model using the original aperture, focus distance, and exposure settings corresponding to each captured view, and compare these synthesized views against the actual captures. This comparison is conducted in LDR space, thus we employ traditional image quality metrics, in particular, PSNR, SSIM, and LPIPS. Quantitative results are presented in Table 6, while qualitative results are illustrated in Fig. 7 for the defocus loss term and in Fig. 8 for the exposure loss term.
@@
Method PU-PSNR $\uparrow$ PU-SSIM $\uparrow$ HDR-VDP $\uparrow$
@@
Ours 34.04 $\pm$ 1.67 0.95 $\pm$ 0.00 9.83 $\pm$ 0.05
Table 6: Evaluation of the impact of the defocus and exposure loss terms in the real dataset.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
4.5 Discussion of Depth-of-Field Generation
Our depth-of-field module can achieve real-time performance at high quality. Alternatives include accumulation buffering [HA90], which delivers high-quality results but is slow due to the need for multiple samples. Another method, used in RawNeRF [MHMB22], decomposes the image into multiplane images at different depths [SGHS98, KS07], applies image space blurring, and then composites them. While this improves speed, it struggles with large depth complexity. In contrast, our approach does not have this limitation.
To shed some light on the respective trade-offs, we compare our method against accumulation buffering for depth of field simulation. Specifically, we replace our DoF module with accumulation buffering, testing with 20 ($\text{AB}_{\text{low}}$) and 120 samples ($\text{AB}_{\text{high}}$). Unlike accumulation buffering, which depends on Monte Carlo sampling, our method requires only a single rendering. As shown in Table  7, our method delivers depth-of-field quality comparable to accumulation buffering with a high sample count while significantly improving speed, enabling real-time rendering. Visual comparisons in Fig. 9 reveal that lower sampling in accumulation buffering results in noticeable artifacts.
Table 7: Comparisons of our depth-of-field rendering module against solutions involving an accumulation buffer.
Method PSNR $\uparrow$ SSIM $\uparrow$ FPS $\uparrow$
$\text{AB}_{\text{high}}$ 27.59 $\pm$ 3.39 0.94 $\pm$ 0.02 1
$\text{AB}_{\text{low}}$ 27.29 $\pm$ 3.34 0.93 $\pm$ 0.02 6
Ours 27.41 $\pm$ 3.30 0.94 $\pm$ 0.02 90
Figure 9: Depth-of-field rendering comparison. Our real-time method achieves high-quality results comparable to those using an accumulation buffer with many samples.
@@ 5 Limitations and Future Work
Our approach does not come without limitations, opening up fruitful avenues for future work. Our defocus model is a two-fold approximation of the actual physical processes in a real camera lens. First, we use a thin-lens model [PC81], which, despite being commonly used, is a simplified representation of defocus. More sophisticated models have been developed in the forward-rendering literature [KMH95] and related work on defocus modeling in the context of reconstruction from image stacks [WSP23b] has shown that advanced effects, such as lens breathing, can significantly improve accuracy. Second, we approximate the thin-lens circle of confusion with a Gaussian function to achieve real-time rendering performance. While this is a commonly used design decision [Fav10, SZW23], it results in a slight softening of bokeh due to the deviation from a disk kernel. We believe that moving towards more realistic lens models has the potential to further increase result quality and cinematic appeal.
In scenes characterized by extremely high dynamic range, where very dark regions may contain significant noise across most input images, artifacts may occur in the reconstruction of these dark areas. These artifacts are pronounced when selecting very high exposures (Fig. 10). This is expected, since there is insufficient useful information, with noise dominating instead in these areas. Nevertheless, these artifacts are noticeable only under extreme exposure settings in the reconstruction, while our method remains capable of delivering pleasing results for the rest of the scene. As discussed above, reduced exposure, whether through shorter exposure times or a smaller aperture, results in increased noise in real camera images. However, this paper primarily focuses on HDR and depth of field characteristics. Addressing noise in dark regions is identified as a potential area for future research. Therefore, for our synthetic dataset, we adhere to the HDR-Nerf [HZF22] procedure to render noise-free images.
@@ 6 Conclusion
Acknowledgements
This work has been partially supported by grant PID2022-141539NB-I00, funded by MICIU/AEI/10.13039/501100011033 and by ERDF, EU.